### v1

@@ Acknowledgements
The first author is a research fellow at Matter Labs. The fourth author is a researcher at Matter Labs. This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.
@@ Appendix 0.A Continuous-time Loss-versus-staking
where we used that $\mathbb{E}[\exp(uZ)]=\exp(\frac{u^{2}}{2})$ for $u\in\mathbb{R}$ and $Z\sim\mathcal{N}(0,1)$.
For the traditional staking we assume a fixed compounding staking rate $r$ that leads to $V_{stake}(t)=V(0)\cdot\exp(rt)$.
The LVS is the difference between the value of the liquidity position at time $t$ and the staking position until time $t$.
For the traditional staking this is therefore given by
$LVS_{stake}(t)=V(0)\exp(rt)-V_{LP}(t)$
and for an LST by
@@
The value of $V_{LP}(t)$ depends on the chosen AMM and corresponding parameters. We discuss two examples in the following sections.
@@ Appendix 0.B LP Losses in CPMMs
$x(t)$ $=\frac{L}{\sqrt{P(t)}}$
$y(t)$ $=L\sqrt{P(t)},$
@@
Loss-versus-staking
The required return to compensate loss-versus-staking is
@@
Loss-versus-holding
@@
And for $\sigma=0$, we have $\mathbb{E}[rr_{LVH}]=(e^{r/2}+e^{-r/2})/2-1$.
@@ Appendix 0.C LP Losses in Concentrated Liquidity CPMMs
Loss-versus-staking
@@
Loss-versus-holding
@@ Appendix 0.D Extra Figures and Tables
wstETH_WETH_100
Uniswap v3
39.66m
58.65m
wstETH_WETH_500
Uniswap v3
0.77m
0.02m
cbETH_WETH_500
Uniswap v3
4.01m
0.7m
rETH_WETH_500
Uniswap v3
4.53m
0.3m
stETH_WETH
Curve v1
30
251.64m
17.72m
stETH_WETH_ng
Curve v1
1500
74.61m
1.25m
stETH_WETH_con
Curve v1
1000
4.35m
0.12m
rETH_WETH
Curve v2
20 000 000
0.01
8.18m
0.6m
cbETH_WETH
Curve v2
20 000 000
0.01
4.29m
0.30m
@@
Figure 8: Historical portfolio value of LPs (blue line), LPs with CRV token rewards (dotted blue line), HOLD (orange line), LST holder (green line), and the average staker (red line) for selected Curve LST pools, denominated in ETH. Pools with a longer history are additionally plotted after the Shanghai upgrade.
Figure 9: Historical 7 day moving average for the selected Curve liquidity pools with LST. Color green marks periods with no loss-versus-staking (LVS) and color yellow periods with no loss-versus-holding (LVH, impermanent loss). Pools with a longer history are additionally plotted after the Shanghai upgrade.
Figure 10: Historical 30 day moving average for the selected Curve liquidity pools with LST. Color green marks periods with no loss-versus-staking (LVS) and color yellow periods with no loss-versus-holding (LVH, impermanent loss). Pools with a longer history are additionally plotted after the Shanghai upgrade.
Figure 11: Historical portfolio value of LPs (blue line), HOLD (orange line),LST holder (green line), and the average staker (red line) for selected Uniswap v3 LST pools, denominated in ETH. Pools with a longer history are additionally plotted after the Shanghai upgrade.
Figure 12: Historical 7 day moving average for the selected Uniswap v3 liquidity pools with LST. Color green marks periods with no loss-versus-staking (LVS) and color yellow periods with no loss-versus-holding (LVH, impermanent loss). Pools with a longer history are additionally plotted after the Shanghai upgrade.
Figure 13: Historical 30 day moving average for the selected Uniswap v3 liquidity pools with LST. Color green marks periods with no loss-versus-staking (LVS) and color yellow periods with no loss-versus-holding (LVH, impermanent loss). Pools with a longer history are additionally plotted after the Shanghai upgrade.

### v2

@@ Acknowledgements
This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use. One of the authors’ research was supported by the EPSRC CDT in Mathematics of Random Systems (EPSRC Grant EP/S023925/1).
@@ Appendix 0.A Continuous-time Loss-versus-staking
where we used that $\mathbb{E}[\exp(uZ)]=\exp(\frac{u^{2}}{2})$ for $u\in\mathbb{R}$ and $Z\sim\mathcal{N}(0,1)$. Hence, the LVS is given by
@@
Here, the value of $V_{LP}(t)$ depends on the chosen AMM and corresponding parameters. We discuss two examples in the following sections.
@@ Appendix 0.B LP Losses in CPMMs
$x(t)=\frac{L}{\sqrt{P(t)}}\qquad\text{and}\qquad y(t)=L\sqrt{P(t)},$
@@
0.B.0.1 Loss-versus-staking.
The required return to compensate LVS is
@@
0.B.0.2 Loss-versus-holding.
The required return to compensate LVH is
@@
And for $\sigma=0$, we have $\mathbb{E}[rr_{LVH}]=(e^{rt/2}+e^{-rt/2})/2-1$.
@@ Appendix 0.C LP Losses in Concentrated Liquidity CPMMs
0.C.0.1 Loss-versus-staking.
@@
0.C.0.2 Loss-versus-holding.
@@ Appendix 0.D Extra Figures and Tables
wstETH_WETH_100 Uniswap v3 - - 39.66m 58.65m
wstETH_WETH_500 Uniswap v3 - - 0.77m 0.02m
cbETH_WETH_500 Uniswap v3 - - 4.01m 0.7m
rETH_WETH_500 Uniswap v3 - - 4.53m 0.3m
stETH_WETH Curve v1 30 - 251.64m 17.72m
stETH_WETH_ng Curve v1 1500 - 74.61m 1.25m
stETH_WETH_con Curve v1 1000 - 4.35m 0.12m
rETH_WETH Curve v2 20 000 000 0.01 8.18m 0.6m
cbETH_WETH Curve v2 20 000 000 0.01 4.29m 0.30m
@@
(a) Curve pools
(b) Uniswap v3 pools
Figure 8: Historical portfolio value of LPs (blue line), LPs with CRV token rewards (dotted blue line), HOLD (orange line), LST holder (green line), and the average staker (red line) for selected Curve and Uniswap v3 LST pools, denominated in ETH. Pools with a long history are additionally plotted after the Shanghai upgrade.
(a) Curve pools
(b) Uniswap v3 pools
Figure 9: Historical 30 day moving average for selected Curve and Uniswap v3 LST liquidity pools. Color green marks periods with no loss-versus-staking (LVS) and color yellow periods with no loss-versus-holding (LVH, impermanent loss). Pools with a longer history are additionally plotted after the Shanghai upgrade.